Find the equation of the tangent plane to the conicoid x^2+y^2 = kz at the point
(k;k;2k), where k is a constant. Represent the plane geometrically. Now take
different values of k, including both positive and negative, and see how the shape of
the conicoid changes.
1
Expert's answer
2019-03-19T13:25:24-0400

We have an equation of conicoid

"F(x,y,z) = x^2 + y^2 - kz"


And the point

"M(k,k,2k)"


Let’s find a derivative

"F\u2019x(x,y,z) = 2x; F'x(M) = 2k"


"F\u2019y(x,y,z) = 2y; F\u2019y(M) = 2k"


"F\u2019z(x,y,z) = -k; F\u2019x(M) = -k"

So equation of tangent plane is:

"F\u2019x(M)*(x-k) + F\u2019y(M)*(y-k) + F\u2019z(M)*(z-2k) = 0"


"2k*(x-k) + 2k*(y-k)*-k*(z-2k) = 2kx + 2ky - kz -2k2 = 0"

Answer:

"2kx + 2ky - kz -2k^2 = 0"

if k > 0




if k < 0